OpenStudy (anonymous):

I need a reminder of how to do these.. I need to rationalize the denominator. 1/sqrtof6

OpenStudy (anonymous):

\[1/\sqrt{6}\]

OpenStudy (lgbasallote):

a reminder hmm...first step is to look at your denominator..make sure it's just one term consist of square root

OpenStudy (anonymous):

Yes, I know it cannot be a square root, but I forget how to turn it into a real number. Do I take the sqrt of 6 and times the whole thing by 2?

OpenStudy (anonymous):

Or do I multiply the sqrt of 6 to the top and bottom?

OpenStudy (lgbasallote):

multiply sqrt 6 to top and bottom

OpenStudy (anonymous):

So it would be \[\sqrt{6}/6?\]

OpenStudy (lgbasallote):

yes

OpenStudy (anonymous):

So that would apply with any sqrt as a denominator? Thank you.

OpenStudy (lgbasallote):

as long as it doesn't look like this \[\frac 1{1 + \sqrt 6}\] if it looks like that, then that's another story